OpenStudy (theprofessor35):

Write the polynomial in standard form: (0.3 - 2m)^2

OpenStudy (theprofessor35):

Someone....help? @mathmale @TSO @Mehek14

OpenStudy (caozeyuan):

(a-b)^2=a^2-2ab+b^2

OpenStudy (theprofessor35):

How'd you find that answer though???

OpenStudy (jdoe0001):

tis a binomial at the second power expand it then put it in standard form

OpenStudy (mathmale):

(0.3 - 2m)^2 is a polynomial; it is NOT an equation. You are asked expand this "square of a binomial" and then rearrange the resulting terms in descending order by powers of x. Try that, please. Do not try to make an equation out of this. Ask questions if any of these suggestions are not clear to you.

OpenStudy (jdoe0001):

or FOIL it \(\bf (0.3 - 2m)^2\iff (0.3 - 2m)(0.3 - 2m)\)

OpenStudy (theprofessor35):

So, what's next? What do I do? I have NO idea what Im being asked here. .-.

OpenStudy (theprofessor35):

@mathmale @jdoe0001

OpenStudy (mathmale):

I would recommend that you do the same thing as jdoe has recommended:\[\bf (0.3 - 2m)^2\iff (0.3 - 2m)(0.3 - 2m)\]

OpenStudy (theprofessor35):

So....that would be my final answer?

OpenStudy (mathmale):

\[(0.3 - 2m)^2\iff (0.3 - 2m)(0.3 - 2m)\]

OpenStudy (mathmale):

No. Here's what's happening here: 1. Copy down the original problem: (0.3-2m)^2. 2. Expand this "square of a binomial" as jdoe has recommended. 3. Write the four resulting terms in descending order by powers of m. 4. Simplify the result from (3), above.

OpenStudy (mathmale):

\[(0.3 - 2m)^2\iff (0.3 - 2m)(0.3 - 2m)\]

OpenStudy (theprofessor35):

So, I still don't get it. My final answer would be 3 after I squared the binomial??

OpenStudy (mathmale):

What does "FOIL" mean? F stands for "first terms," O for "outer terms," I for "inner terms," and L for "last terms." \[ (0.3 - 2m)(0.3 - 2m)\]

OpenStudy (mathmale):

F: First terms: 0.3 and 0.3 are the first terms in the above product of 2 binomials. Multiply them together. What do you get? O; Outer terms. 0.3 and -2m are the outer terms of the above product. Multiply them together. What do you get? Can y ou do the I and L parts on your own? I: innter terms; L: Last terms.

OpenStudy (mathmale):

(0.3)(0.3)=? (0.3)(-2m)=? (-2m)(0.3)=? (-2m)(-2m)=?

OpenStudy (theprofessor35):

0.3 * 0.3 = 0.09 0.3*-2m = -0.6m -2m)(0.3 = -0.6m -2m*-2m = 4m

OpenStudy (mathmale):

Then sum them up: 0.09 - 0.6m - 0.6m + 4m^2 (note that you must square that m!) OK with this?

OpenStudy (theprofessor35):

Ohhhh, thank soooooooooooooooooooooooooooooooooooooo much :)

OpenStudy (mathmale):

Combine the 2 middle terms into one. Then, reverse the order of these terms, so that the first term is the m^2 term, the 2nd is the m term and the 3rd is the constant 0.09. Then you'll be done. You're welcome. Glad to work with you.

OpenStudy (theprofessor35):

Wait what? What do you mean by "reverse"? Sorry, I don't know alot of math, math just isn't my best subject. :( So, you meant that I do this:- 0.09 - 0.6m - 0.6m + 4m^2 4m^2 + 0.6m - 0.6m - 0.09

OpenStudy (theprofessor35):

Or did you mean to combine like terms?

OpenStudy (mathmale):

Reverse order, as in 4 3 2 1 becomes 1 2 3 4. Yes, that's what I meant. However, you still have to combine the 2 middle terms, and to realize that the last (the constant) term is +, not -.

OpenStudy (theprofessor35):

So, my final answer would be: 4m^2 + m + 0.09

OpenStudy (theprofessor35):

Am I right? :)

OpenStudy (theprofessor35):

@mathmale

OpenStudy (mathmale):

What's the coefficient of m? what is -0.6m - 0.6m?

OpenStudy (theprofessor35):

Oh, -0.6 + -0.6 = -0.12 So, it would be: -0.12m

OpenStudy (theprofessor35):

-1.2m

OpenStudy (theprofessor35):

correction, sorry

OpenStudy (mathmale):

You are to combine these two terms, both of which are -. close, but not quite. Please add: -0.6m -0.6m -------

OpenStudy (mathmale):

Final answer: for the whole polynomial:

OpenStudy (theprofessor35):

4m^2 -1.2m + 0.09

OpenStudy (mathmale):

Great.